OpenStudy (anonymous):

Dosha is creating a new dessert that has two layers shaped like cones. The inner cone is frozen ice cream and has a diameter of 12 cm and a height of 6 cm. The outer layer is a thin wafer shell, like an upside-down ice cream cone, with a height of 15 cm and the same diameter as the inner layer. Dosha will inject a cream filling into the space. What is the volume of the cream filling? Use 3.14 to approximate pi and express your final answer in hundredths. ______ cm3

OpenStudy (anonymous):

@perl @paki ?? can you help me with this

OpenStudy (anonymous):

thanks a lot for the help

OpenStudy (perl):

so we have two cones?

OpenStudy (anonymous):

yh

OpenStudy (perl):



OpenStudy (perl):



OpenStudy (anonymous):

mhm

OpenStudy (perl):

Volume of filling = Volume of Total cone - Volume of inner cone = 1/3 * Pi * 6^2 * 15 - 1/3 * Pi * 6^2 * 6

OpenStudy (anonymous):

ok let me do it

OpenStudy (perl):

\[\large Volume ~of~ cone= \frac{1}{3}\pi r^2 h\]

OpenStudy (perl):

oh i notice it says use 3.14 for Pi

OpenStudy (perl):

Volume of filling = Volume of Total cone - Volume of inner cone = 1/3 *3.14 * 6^2 * 15 - 1/3 * 3.14 * 6^2 * 6

OpenStudy (anonymous):

339.12.... sorry for the late response